OpenStudy (shamil98):

Let f and g be the functions given by f(x) = 2x(1-x) and g(x) = 3(x-1)√x for 0 ≤ x ≤ 1. Find the volume of the solid generated when the shaded area enclosed by the graphs of f and g is revolved about the horizontal line y = 2. I used the washer method for this, and input it into my graphing calc and got -6.48.. it's not supposed to be negative so i did something wrong.

OpenStudy (shamil98):

The graph of the functions:

OpenStudy (shamil98):

sec, won't let me post attached file..

OpenStudy (shamil98):

http://i.imgur.com/oRlBJlw.png this thingy

OpenStudy (anonymous):

looks like washer method

OpenStudy (nincompoop):

yeah it says that he needs to use the washer method

OpenStudy (anonymous):

We need your work to find the mistake

OpenStudy (nincompoop):

nevermind, he said he used the washer LAUGHING OUT LOUD

OpenStudy (anonymous):

\[ \pi\int\limits_0^1[g(x)]^2-[f(x)]^2\;dx \]

OpenStudy (anonymous):

well, not quite.

OpenStudy (anonymous):

Gotta put in that 2-g / 2-f

OpenStudy (shamil98):

Oh, I did that but reversed the order.. i put (2-f(x))^2 - (2-g(x))^2 dx and what i got is -19.113, forgot to multiply the pi after integrating.

OpenStudy (anonymous):

\[ \pi\int\limits_0^1[2-f(x)]^2 - [2-g(x)]^2\;dx \]

OpenStudy (shamil98):

wait what.

OpenStudy (anonymous):

basically, it's (outer radius)^2 - (inner radius)^2

OpenStudy (shamil98):

I did that already wio, i got the -19.113. is it supposed to be negative then?..

OpenStudy (anonymous):

There is no way for the integrand to be negative.

OpenStudy (anonymous):

probably you just made some algebraic mistake

OpenStudy (shamil98):

Oh i was putting this in.. \[{\pi}\int\limits_{0}^{1} (2-(2x(x-1))^2 - (2-(3(x-1)\sqrt{x}))^2 dx\]

OpenStudy (nincompoop):

use square brackets and {} also so it is less confusing